Knockout3000:

The first sequence rule is multiply by 3 starting from 4. The second sequence rule is add 8 starting from 20. What is the first number that appears in both sequences? 12 20 28 36

Knockout3000:

help?

AquarytheALPHA:

ill help

Knockout3000:

@aquarythealpha wrote:
ill help
Nvm someone just dm'd me the answer

AquarytheALPHA:

oop

AquarytheALPHA:

well i tried

Knockout3000:

it's ok. i will give you fan and medal.

AquarytheALPHA:

ty

Knockout3000:

Rule 1: Multiply by 2, then add one third starting from 1. Rule 2: Add one half, then multiply by 4 starting from 0. What is the fourth ordered pair using the two sequences? (5, 10) (four and two thirds, 42) (21, 170) (two and one third, 2)

Knockout3000:

I'M NOW RED!!!!!!!!!!!!!!!!!!!!!!!!! 

AquarytheALPHA:

bruh

AquarytheALPHA:

it could be B or D

Knockout3000:

Which one

Knockout3000:

question

AquarytheALPHA:

the second one

Knockout3000:

This one?

@knockout3000 wrote:
Rule 1: Multiply by 2, then add one third starting from 1. Rule 2: Add one half, then multiply by 4 starting from 0. What is the fourth ordered pair using the two sequences? (5, 10) (four and two thirds, 42) (21, 170) (two and one third, 2)

AquarytheALPHA:

@knockout3000 wrote:
This one?
@knockout3000 wrote:
Rule 1: Multiply by 2, then add one third starting from 1. Rule 2: Add one half, then multiply by 4 starting from 0. What is the fourth ordered pair using the two sequences? (5, 10) (four and two thirds, 42) (21, 170) (two and one third, 2)
Yesss

Knockout3000:

@candyplays321 dm'd the answer.

surjithayer:

for first \[y=4(3)^{n-1}\] for second \[y=20+8(n-1)\] for n=1 \[4(3)^{1-1}=20+8(1-1)\] 4=20 for n=2 \[4(3)^{2-1}=20+8(2-1)\] \[4(3)=20+8\] 12=28 not true for n=3 \[4(3)^{3-1}=20+8(3-1)\] \[4(3)^2=20+8(2)\] 36=36 so third term =36 and first term equal=36